OpenStudy (anonymous):

Graph the following linear inequality: 2x+y>4

OpenStudy (primeralph):

y>4-2x. Plot y=4-2x and mark the region less than y.

OpenStudy (anonymous):

How do I plot it? (-2,0) & (0,4)?

OpenStudy (primeralph):

@iambatman Can you explain?

OpenStudy (anonymous):

@adrynicoleb @SolomonZelman , can you help me please?

hero (hero):

1. Isolate y: y > 4 - 2x 2. Plot y = 4 - 2x as a dotted line on an xy coordinate plane 3. Shade everything above the dotted line 4. Write "2x + y > 4" on your graph

OpenStudy (anonymous):

I still don't understand...what points do I need to plot?

hero (hero):

y = 4 - 2x is a "line" that you plot. You only need two points to plot a line. It is best to find the x and y intercepts. To find the x-intercept, let y = 0, then solve for x: 0 = 4 - 2x 2x = 4 x = 4/2 x = 2 So the first point you will plot is (2,0) To find the y-intercept, let x = 0, then solve for y: y = 4 - 2(0) y = 4 - 0 y = 4 So the second point you will plot is (0,4) After plotting both points draw a dotted line through them.